OpenStudy (anonymous):

if y= (3)/(4+x^2), then (dy)/(dx)= if y= (3)/(4+x^2), then (dy)/(dx)= @Calculus1

OpenStudy (anonymous):

y= (3)/(4+x^2) y=(3)(4+x^2)^-1 dy/dx=3(-1)(4+x^2)^-2 (2x) =-6x/(4+x^2)^2 ans